1. Attendance at Orlando’s newest Disneylike attraction, Lego World, has been as follow:


QUARTER

GUEST (IN THOUSADS)

QUARTER

GUEST (IN THOUSADS)

Winter Year 1

73

Summer Year 2

124

Spring Year 1

104

Fall Year 2

52

Summer Year 1

168

Winter Year 3

89

Fall Year 1

74

Spring Year 3

146

Winter Year 2

65

Summer Year 3

205

Spring Year 2

82

Fall Year 3

98


a) Compute the Seasonal Indices using all the data.

b) If the expected guest for the next year (i.e., Forth Year) is 550000, forecast the quarterly attendance for the year 4?






1
Expert's answer
2021-01-04T20:38:14-0500

a)


"Guests\ (Demand) \\"


"\begin{matrix}
 
Quarter & Year1& Year2 & Year3 \\
 Winter & 73 & 65 & 89 \\
Spring & 104 & 82 & 146 \\
Summer & 168 & 124 & 205\\
Fall & 74 & 52 & 98
\end{matrix}"

"Average\ Demand=\dfrac{73+65+89}{3}=\dfrac{227}{3}=75.67"

"Average\ Demand=\dfrac{104+82+146}{3}=\dfrac{332}{3}=110.67"

"Average\ Demand=\dfrac{168+124+205}{3}=\dfrac{497}{3}=165.67"

"Average\ Demand=\dfrac{74+52+98}{3}=\dfrac{224}{3}=74.67"

"Total \ Average \ annual\ demand"

"=\dfrac{227+332+497+224}{3}=\dfrac{1280}{3}=426.67"

"Total \ Average \ quaterly\ demand"

"\dfrac{1280}{3(4)}=\dfrac{320}{3}=106.67"


"Seasonal \ index =\dfrac{\dfrac{227}{3}}{\dfrac{320}{3}}=\dfrac{227}{320}=0.709375"

"Seasonal \ index =\dfrac{\dfrac{332}{3}}{\dfrac{320}{3}}=\dfrac{332}{320}=1.037500"

"Seasonal \ index =\dfrac{\dfrac{497}{3}}{\dfrac{320}{3}}=\dfrac{497}{320}=1.553125"

"Seasonal \ index =\dfrac{\dfrac{224}{3}}{\dfrac{320}{3}}=\dfrac{224}{320}=0.700000"

"\begin{matrix}
 
Quarter & Seasonal \ index\\
 Winter & 0.709375 \\
Spring & 1.037500 \\
Summer & 1.553125 \\
Fall & 0.700000
\end{matrix}"

b)


"Winter=\dfrac{550}{4}\times 0.709375=97.539 (thousands)"

"Spring=\dfrac{550}{4}\times 1.037500=142.656(thousands)"

"Summer=\dfrac{550}{4}\times 1.553125=213.555 (thousands)"

"Fall=\dfrac{550}{4}\times 0.700000=96.250 (thousands)"


"\begin{matrix}
 
Quarter & Seasonal \ index & Guest (thousands)\\
 Winter & 0.709375 & 97.539 \\
Spring & 1.037500 & 142.656 \\
Summer & 1.553125 & 213.555\\
Fall & 0.700000 & 96.250
\end{matrix}"